Given a nominnal interest rate of 6 percent in which of the following casses would you earn the lowest after tax real interest rate
1
Expert's answer
2020-04-20T18:55:41-0400

Nominal Return x (1 - tax rate)=0.06(1-0.05)=0.057

Nominal Return x (1 - tax rate)=0.06(1-0.2)=0.048

Nominal Return x (1 - tax rate)=0.06(1-0.3)=0.042


In the third option, the lowest rate after tax. The higher the tax rate, the lower the nominal rate